chloe2:

Which of the following is not a subset of the real number Natural numbers Whole numbers Rational numbers none of the above

Narad:

\[\mathbb{N} \subset \mathbb{Z}^+ \subset \mathbb{Q} \subset \mathbb{R} \]

chloe2:

which one is it